OpenStudy (anonymous):

Solve 3x+y=-3 , y=x+5 Solve by using substitution. Express your answer as an ordered pair Please steb by step

OpenStudy (anonymous):

3x+(x+5)=-3 4x=-3-5 4x=-8 x=-2 substitute to y=x+5 y=3